# Towards Active Synthetic Data Generation for Finetuning Language Models

### v1

@@ 5.4 Results
Using synthetic data yields significant gains in student capabilities compared to a fixed seed dataset of the same size (Section 5.4.1), additionally iterative synthetic data generation is more data efficient than static generation (Section 5.4.2). Note that static generation is equivalent to random sampling of prompts for synthetic data generation in our setting, since it is not conditioned on the current state of the student. Finally, we analyse our synthetically generated data to show that it retains important properties of the original seed data, (Section 5.4.3). Unless stated otherwise, results are a mean and standard deviation over $3$ independent runs.
5.4.1 Training on synthetic data improves performance
SFT on synthetic data results in significantly improved capabilities when compared to using the original seed dataset.  Figure 2 compares SFT performance on the seed data to synthetic data of equal size, showing a dramatic increase in performance across all datasets when doing SFT on synthetic question-answers pairs. In the same figure, we see large increases in performance when using synthetic answers $z_{i}=(x_{i},\hat{y}_{i})$ instead of seed answers $y$, likely due to better formatting and high quality chain-of-thought in synthetic answers. In Game of 24 there is a small drop in performance when training on synthetic questions and answers compared to synthetic answers only, showing that the generation of novel questions by the teacher yields some lower quality synthetic questions. Regardless, next we show how this enables us to scale dataset
@@ 5.4.1 Training on synthetic data improves performance
sizes efficiently.
5.4.2 Iterative Generation Yields More Data-Efficient Results
Active selection is more data efficient than random sampling for generating productive synthetic data, resulting in better performance using fewer samples. In Figure 3, we can see that random sampling underperforms when compared to certain active selection methods across all datasets. Among these, we find that simply prioritizing high-loss data consistently performs well.
@@ 5.4.2 Iterative Generation Yields More Data-Efficient Results
To compare algorithms across all datasets we can aggregate results as a pairwise winrate matrix $\bm{P}$}. We increment ${\bm{P}}_{ij}$ if $\bm{1}\{\hat{\mu}_{i}-\alpha\cdot\hat{\text{se}}_{i}>\hat{\mu}_{j}+\alpha\cdot\hat{\text{se}}_{j}\}$, where $\hat{\mu}_{i}$ is the sample mean and $\hat{\text{se}}_{i}$ is the standard error of the performance of algorithm $i$ for a dataset, for a particular dataset size, and $\alpha$ is the confidence level which we set to $1$ (making it a $68\%$ confidence interval). By summing the “wins” across the rows and normalizing we can understand how often algorithms are outperformed on average. Column-wise averages are shown in the last row, where lower is better, to understand which algorithm is more data efficient in total. We find that using random sampling is outperformed by various other methods that use the student model to guide synthetic data generation (Figure 4).
@@
We can glean from Figure 4 that the best selection algorithm simply uses a high loss to select data that are difficult for the SLM. LLM-as-a-judge also tends to be effective, though by a smaller margin. Interestingly, BADGE and Lion, which both aim to select diverse data, do not perform much better than random sampling (Figure 4). It is worth mentioning that, because of the need to access a teacher model for scoring, LLM-as-a-judge is computationally demanding. Assuming that the cost of evaluating the teacher model dominates the cost of evaluating the student, a common assumption in the active learning literature and a reasonable assumption as the number of parameters of the teacher model can be $3$ orders of magnitude larger than the student models we consider. Then, if we consider the total number of teacher input and output tokens as a budget instead of the number of generated samples, Lion and LLM-as-a-judge (hard) are far more expensive than other methods (Figure 5). This additional compute is better allocated towards simply generating more synthetic data with a cheaper and more effective selection strategy.
Reward scoring also requires an external model, but because we can use a reward model that has the same number of parameters as our student, calls to the reward model are generally less expensive than to a teacher—we opt to not treat them in the same way and do not count the number of input tokens to the reward model in Figure 5. Overall random selection requires between $33\%$ to $100\%$ more SFT data to obtain the same performance as the best selection methods across all datasets (Figure 3). For $2/4$ of these datasets iterative synthetic data generation using the loss on the student’s own predictions leads to more data efficient results compared to prior works that perform SFT using similarly sized datasets on the same SLM,
@@
Figure 5: Iterative synthetic data generation learning curves on GSM8k: student performance versus the number of teacher input and output tokens. The total number of input and output tokens are a proxy for the amount of compute used by the teacher for various selection methods.
see Section C.2 for comparisons.
5.4.3 On the fidelity of synthetic data to its original data
Original dataset scores are retained by synthetic data. When we measure the median score over the entire selected dataset and the corresponding median over the synthetic dataset we find a very high rank correlation over the course of iterative synthetic data generation. This shows that at a dataset level the order of scores—uncertainty (measured by the loss) or data quality (measured by the reward)—from the iterative synthetic dataset curriculum are preserved (Figure 6). In spite of this, synthetic data generation is an uncontrollable and noisy process which perturbs data by rephrasing, complicating or simplifying and adding chain-of-thought rationales. This is manifested when we plot the scores of the individual data points before and after synthetic data generation (Figure 6, centre). We notice that the individual data point rank correlations are low but significantly greater than zero across all datasets (Section C.3).
@@ 5.4.3 On the fidelity of synthetic data to its original data
Figure 6: The rank correlations between original and synthetic dataset scores from iterative synthetic data generation. We plot student loss and reward scores and show Spearman’s rank correlations ($\rho$) between dataset medians before and after synthetic data generation. We zoom in on relationships at an individual data-point level where there is low correlation between the original and synthetic data point scores (centre). The red line is the line of best fit to the data. All rank correlations are highly significant with a p-value of $p<0.001$.
Prioritizing “difficult” data results in more difficult synthetic data than random sampling. We compare the cumulative accuracy of synthetic data ordered randomly minus the cumulative accuracy of synthetic data ordered with a high to low original loss and low to high reward in Figure 7. For GSM8k and Math1-3 the difference in the cumulative accuracies is significantly greater than 0. So prioritizing data according to these scores results synthetic data with lower student accuracies than random sampling. The synthetic data is “harder” using these active learning approaches and these “hardness” qualities are integrated into the synthetic data the teacher generates. This is also seen for the first iteration for the Game of 24 dataset. In contrast, at the final iteration of the Game of 24 dataset and for the ProntoQA dataset the student obtains high accuracies on the synthetic data so we cannot see any difference in cumulative accuracies. Using low reward selection produces more difficult samples compared to random for the Game of 24 dataset (Figure 7), and results in poorer SFT performance than random however (Figure 3). This is because we observe that the reward scorer introduces other biases; it is biased to long student answers and so biased toward a narrow stylistic set of student responses, as observed in other work (Shen et al., 2023; Bu et al., 2025). Moreover each selection algorithm has its own selection biases which results in different synthetic dataset token distributions, see Section C.5. See Figure 11 for the original cumulative accuracies used for calculating the percentage differences in Figure 7.
@@
Figure 7: The percentage difference in synthetic data cumulative accuracy when using random sampling minus score ordering: high to low loss and low to high reward. For each original data point we score it using the student model from the first and final iteration of iterative synthetic data generation (rows). See Figure 11 for the cumulative accuracies for individual replicates.
@@ 6 Conclusion and Discussion
Synthetic data are extremely effective for finetuning SLMs, enabling substantial capability improvements. In this work, we focus on supervised finetuning with synthetic data. We demonstrate that iterative synthetic data generation is the most effective strategy for fine-tuning SLMs under a fixed training data budget. By adapting teacher generation to the evolving state of the student model, this approach creates a natural curriculum that consistently outperforms static synthetic datasets in both performance and data efficiency. Furthermore, in line with Occam’s razor, we find that simple data selection methods, such as prioritizing hard samples with high loss, outperform complicated and expensive LLM-as-a-judge based methods. We also find that the student can steer the synthetic data generation since firstly dataset level statistics are preserved by synthetic data generation. Secondly, prioritizing hard samples in turn produces hard synthetic samples. As a result, the synthetic data retains important characteristics of the student selected data, enabling enhanced student capabilities after finetuning. Limitations are overviewed in Appendix A.
@@ Appendix D Dataset further details
We perform SFT on a Mistral-7B-Instruct-v0.3 [Jiang et al., 2023a] student on school level mathematics questions [Cobbe et al., 2021]. We use an external language model gpt4o-mini to assess whether the student’s answer is equivalent to the ground truth answer, in a similar manner to Mitra et al. [2024], see Section D.3 for prompting details. We take $748$ question-answer pairs from the test set as a validation set and use $500$ question-answer pairs as a test set***https://huggingface.co/datasets/openai/gsm8k.
@@
The questions are synthetically generated logical chain-of-thought style reasoning questions with boolean answers [Saparov and He, 2023]. We perform SFT with a Qwen1.5-7B-Chat student model. We use an external language model gpt4o-mini to assess whether the student’s reasoning steps are correct and whether the student answer is equivalent to the ground truth answer like for GSM8k, see Section D.3 for details. We use $300$ question-answer pairs as a validation set and the remaining $200$ as a test setWe usehttps://huggingface.co/datasets/renma/ProntoQA for validation and testing, as a train set we use https://huggingface.co/datasets/longface/prontoqa-train like in [Huang et al., 2024], questions are distinct between these two ProntoQA datasets..

### v2

@@ 5.4 Results
This subsection presents our main results, which includes (1) performance comparisons between fitting the SLM on seed data and synthetic data (Section 5.4.1), and (2) between standard active learning strategies and more modern, LLM-based alternatives (Section 5.4.2). Further, we (3) analyse synthetically generated data to demonstrate it retains important properties of the original seed data, providing the underlying property of this mechanism affording the effectiveness of active methods (Section 5.4.3). Throughout this section, note that static generation is equivalent to random sampling of prompts for synthetic data generation in our setting, as it is not conditioned on the current state of the student. Unless stated otherwise, results show the mean and standard deviation over $3$ independent runs***Website and code: https://iterative-sd.github.io/.
5.4.1 Synthetic Data Improves Performance
SFT on synthetic data results in significantly improved capabilities when compared to using the original seed dataset.  Figure 2 compares SFT performance on the seed data to synthetic data of equal size, showing a dramatic increase in performance across all datasets when doing SFT on synthetic question-answers pairs. In the same figure, we see large increases in performance when using synthetic answers $z_{i}=(x_{i},\hat{y}_{i})$ instead of seed answers $y$, likely due to better formatting and high quality chain-of-thought in synthetic answers. In Game of 24 there is a small drop in
@@ 5.4.1 Synthetic Data Improves Performance
performance when training on synthetic questions and answers compared to synthetic answers only, showing that the generation of novel questions by the teacher yields some lower quality synthetic questions. Regardless, next we show how this enables us to scale dataset sizes efficiently.
5.4.2 Iterative Generation is More Data Efficient than Static Generation
Active selection is more data efficient than random sampling for generating productive synthetic data, resulting in better performance using fewer samples. Figure 3 shows learning curves, with each plot measuring the test accuracy of a given selection method as a function of the labelling budget; each point is an active learning iteration. As mentioned earlier, random selection, because it is not conditioned on the current state of the student, is equivalent to the typical approach of static generation at the indicated data size. We find that this technique is often outperformed by a student-in-the-loop alternative—horizontal lines indicate the number of additional samples that would be required by static, random sampling in order to achieve the same performance as the best active learner in the corresponding plot (between $1.3\times$ and $2\times$). We find that uncertainty sampling performs roughly as well, and often better than LLM-based scoring methods. For some datasets, like Game of 24, reward-based methods do quite poorly, likely as a result of the task being out of distribution for the reward model.
Figure 4 aggregates performance differences between all selection strategies and model-dataset pairs considered in this paper. Each experiment composing this figure produced a learning curve in Figure 3, with each method producing a different SLM test accuracy for a variety of generation budgets. Here, we aggregate results by measuring which algorithms outperform their peers at each generation budget across all models and datasets.
@@ 5.4.2 Iterative Generation is More Data Efficient than Static Generation
Specifically, we aggregate results as a pairwise winrate matrix $\bm{P}$}. We increment ${\bm{P}}_{ij}$ if $\bm{1}\{\hat{\mu}_{i}-\alpha\cdot\hat{\text{se}}_{i}>\hat{\mu}_{j}+\alpha\cdot\hat{\text{se}}_{j}\}$, where $\hat{\mu}_{i}$ is the sample mean and $\hat{\text{se}}_{i}$ is the standard error of the performance of algorithm $i$ for a dataset, for a particular dataset size, and $\alpha$ is the confidence level which we set to $1$ (making it a $68\%$ confidence interval). By summing the “wins” across the rows and normalizing we can understand how often algorithms are outperformed on average. Column-wise averages are shown in the last row, where lower is better, to understand which algorithm is more data efficient in total. We find that random sampling—equivalent to static generation of data—is outperformed by various other methods that use the student model to guide synthetic data generation (Figure 4).
We can glean from Figure 4 that the highest-performing approach is simply uncertainty sampling, using the SLM’s loss on its own generations. LLM-as-a-judge also tends to be somewhat effective, though by a reduced margin. Interestingly, BADGE and Lion, which both aim to select diverse data, do not perform much better than random sampling (Figure 4). This is likely because synthetic data generation is noisy (Section 5.4.3), which diverse selection may exacerbate.
@@
Because of the need to access a teacher model for scoring, LLM-as-a-judge is computationally demanding. Assuming that the cost of evaluating the teacher model dominates the cost of evaluating the student, a common assumption in the active learning literature and a reasonable assumption as the number of parameters of the teacher model can be $3$ orders of magnitude larger than the student models we consider. Then, if we consider the total number of teacher input and output tokens as a budget instead of the number of generated samples, Lion and LLM-as-a-judge (hard) are
@@
Figure 5: Active learning curves on GSM8k: student performance against the number of teacher input and output tokens. The total number of input and output tokens are a proxy for the amount of compute used by the teacher for various methods.
far more expensive than other methods (Figure 5). Our results suggest this additional compute is better allocated towards simply generating more synthetic data with a cheaper and more effective selection strategy, like uncertainty sampling.
@@
Figure 6: The rank correlations between original and synthetic dataset scores from iterative synthetic data generation. We plot student loss and reward scores and show Spearman’s rank correlations ($\rho$) between dataset medians before and after synthetic data generation. We zoom in on relationships at an individual data-point level where there is low correlation between the original and synthetic data point scores (centre). The red line is the line of best fit to the data. All rank correlations are highly significant ($p<0.001$).
Reward scoring also requires an external model, but because we can use a reward model that has the same number of parameters as our student, calls to the reward model are generally less expensive than to a teacher—we opt to not treat them in the same way and do not count the number of input tokens to the reward model in Figure 5. Overall random selection requires between $33\%$ to $100\%$ more SFT data to obtain the same performance as the best selection methods across all datasets (Figure 3). For $2/4$ of these datasets, iterative synthetic data generation using the loss on the student’s own predictions leads to more data efficient results compared to prior works that perform SFT using similarly sized datasets (see Section C.2 for comparisons).
5.4.3 Fidelity of Synthetic Data to its Original Data
Synthetic data generation is a noisy process, perturbing data by rephrasing, complicating or simplifying and adding chain-of-thought rationales. As such the score—either uncertainty (measured by the loss) or data quality (measured by the reward)—of an individual seed sample and its corresponding generated sample appear to have little to do with each other. In aggregate, however, we find high rank correlation between the median score of the seed and generated datasets (Figure 6). This relationship is the underlying principal governing why careful selection of the seed question-response pairs is important: Generated samples inherit underlying properties from the data used to produce them. These attributes, such as the SLM loss, shape the student’s optimization trajectory and generalization capabilities.
@@ 5.4.3 Fidelity of Synthetic Data to its Original Data
Figure 7: The percentage difference in synthetic data cumulative accuracy between samples ordered by score and randomly shuffled. Data are sorted either by uncertainty (high to low) or reward (low to high). Positive values suggest that score ordering picks more difficult synthetic samples in turn yielding lower accuracies. For each original data point we score it using the student model from the first and final iteration of iterative synthetic data generation (rows). See Figure 11 for the cumulative accuracies for individual replicates.
Unsurprisingly, samples with high uncertainty, again measured using the student’s loss on its own generations, tend to also be samples for which the student model struggles to obtain a correct answer. Figure 7 sorts samples by this uncertainty and compares the model’s accuracy on these data in comparison to a random shuffle, this is equivalent to random selection. Specifically, we plot the cumulative percent difference between the SLM accuracy on samples sorted by loss and the SLM accuracy on randomly ordered data. The curve in Figure 7 presents this as a function of the number of samples being used in the calculation, and is repeated for scores from a reward model. In both cases, prioritizing data in this fashion often effectively prioritizes low-accuracy samples, as indicated by the curve’s positive values. The trend is least clear for the ProntoQA dataset, which shows reward scoring as positive and uncertainty scoring as neutral or negative—unlike for other datasets, reward scoring was indeed a slightly more effective selection strategy (fig. 3). Low reward selection also selects lower-accuracy samples compared to random for the Game of 24 dataset (Figure 7), but still performs more poorly in terms of SFT performance than random (Figure 3). This is because the reward scorer introduces biases; specifically, it prefers longer responses, an effect that has been observed in other work (Shen et al., 2023; Bu et al., 2025).
@@ 6 Conclusion and Discussion
This work shows that student-in-the-loop synthetic data generation yields more data-efficient SLM improvements than static one-shot generation, and that simple, active-learning–inspired criteria for selecting seed examples outperform more elaborate LLM-based judging. We demonstrate that synthetic data are partially steerable, with teacher outputs reflecting the properties of selected seeds. These results highlight the value—and underexploration—of principled, effective selection strategies for advancing SLM training. Limitations are overviewed in Appendix A.
@@ Appendix D Dataset Further Details
We perform SFT on a Mistral-7B-Instruct-v0.3 [Jiang et al., 2023a] student on school level mathematics questions [Cobbe et al., 2021]. We use an external language model gpt4o-mini to assess whether the student’s answer is equivalent to the ground truth answer, in a similar manner to Mitra et al. [2024], see Section D.3 for prompting details. We take $748$ question-answer pairs from the test set as a validation set and use $500$ question-answer pairs as a test sethttps://huggingface.co/datasets/openai/gsm8k.
@@
The questions are synthetically generated logical chain-of-thought style reasoning questions with boolean answers [Saparov and He, 2023]. We perform SFT with a Qwen1.5-7B-Chat student model. We use an external language model gpt4o-mini to assess whether the student’s reasoning steps are correct and whether the student answer is equivalent to the ground truth answer like for GSM8k, see Section D.3 for details. We use $300$ question-answer pairs as a validation set and the remaining $200$ as a test set§§§We usehttps://huggingface.co/datasets/renma/ProntoQA for validation and testing, as a train set we use https://huggingface.co/datasets/longface/prontoqa-train like in [Huang et al., 2024], questions are distinct between these two ProntoQA datasets..